Using the information in the table below, calculate the case-mix index for each month in the first quarter at Community Hospital. Round to four decimal places.

Community Hospital

First Quarter Statistics, 20XX

Case-Mix Report

January

February

March

Total discharges

3,213

2,456

3,167

Total weight

7245.2785

7458.7499

8917.4576

Case-mix index
Answer #1

THE COST MIX INDEX IS THE INDEX WHICH IS USED IN HEALTH CARE FACILITY TO REFLECT THE COST OF MEDICARE AND THE TREATMENT REQUIRED BY AN INPATIENT GROUP.

IN SIMPLER TERM THE HIGHER THE CASE MIX INDEX THE MORE SICKER THE INDIVIDUAL IS AND ALSO HE NEEDS MORE COST TO BE SPENT OVER HIS TREATMENT, THE FINANCE DEPARTMENT OF THE HOSPITAL KEEPS THE DATA FOR THEIR ANNUAL EXPENDITURE ON THE PATIENT IN ORDER TO MAINTAIN THE RECORD AND ALSO FOR REIMBURSEMENT.

CMI is calculated by adding all the DRG weight and dividing it by a number of the patient for a specific period.

in most of the hospital cost mix index is calculated annually or monthly.

each DRG is listed during the treatment period along with its weight and it is divided by the number of total individuals DRGs.

in the above question, the data given is on the basis of monthly expenditure.

so,the above

JANUARY COST MIX INDEX = TOATL WEIGHT DRGS / TOATAL NO OF INDIVIDUAL DRGs

= 7245.2785 / 3213

CMI = 2.2549

FEBRUARY COST MIX INDEX   = TOATL WEIGHT DRGS / TOATAL NO OF INDIVIDUAL DRGs

= 7458.7499 / 2456

CMI = 3.0369

MARCH COST MIX INDEX = TOATL WEIGHT DRGS / TOATAL NO OF INDIVIDUAL DRGs

= 8917.4576 / 3167

CMI = 2.8257

THE HIGHER CMI IS IN THE MONTH OF FEBRUARY SO, THE COST OF TREATMENT WILL BE HIGHER FOR THE MONTH OF FEBRUARY, AS HIGHER THE CMI MORE THE SICKER INDIVIDUAL IS.